OpenStudy (anonymous):

Classify the triangle based on the side lengths 9, 40 and 41.

Parth (parthkohli):

What's a triangle called if it has all sides unequal?

OpenStudy (anonymous):

no triangle can be formed with given side lengths, maybe? or its OBtuse?

Parth (parthkohli):

Plus, you may see if it is right, obtuse or acute with their respective formulas.

OpenStudy (lgbasallote):

you can form a triangle here

Parth (parthkohli):

\(a^2 + b^2 = c^2\) Right ^ \(a^2 + b^2 > c^2\) Acute ^ \(a^2 + b^2 < c^2\) Obtuse ^ c is the longest side.

OpenStudy (anonymous):

so the only answer is Acute, because its all angles are less than 90?

OpenStudy (lgbasallote):

you're not given angles

Parth (parthkohli):

They are the sides.

Parth (parthkohli):

You compare \(9^2 + 40^2\) and \(41^2\) and check what's greater.

OpenStudy (anonymous):

so its A, since B, C, D are given angels?

Parth (parthkohli):

Umm. I think you are confused here. Just compare \(9^2 + 40^2\) and \(41^2\).

OpenStudy (anonymous):

Yes I am..

Parth (parthkohli):

What's \(9^2 + 40^2\)?

OpenStudy (anonymous):

81+1600= 1681

Parth (parthkohli):

Correct. Now what's \(41^2\)?

OpenStudy (anonymous):

woah.. same asnwer... so its a RIGHT triangle? B?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

Thanks @ParthKohli

Parth (parthkohli):

You're welcome!